# More variable circadian rhythms in epilepsy: a retrospective cross-sectional study using long-term heart rate recordings from wearable sensors

### v2

@@ Abstract
Background: The circadian rhythm aligns physiology and behaviour with the 24-hour light-dark cycle, and its disruption is linked to neurological disorders such as epilepsy. Disruptions can manifest in different ways, and as such the optimal way to investigate circadian disruption is unclear. A promising but under-explored approach is to assess the intra-individual variability in circadian rhythms over timescales of weeks to years, which is yet to be studied in epilepsy.
Methods: We retrospectively used over 70,000 hours worth of wearable smartwatch data (Fitbit) from 143 people with epilepsy (PWE) and 31 healthy controls. For each participant, we extracted the circadian oscillation underlying their heart rate time series and analysed the intra-individual variability of three circadian properties: period, acrophase, and amplitude.
Findings: We found increased intra-individual variability in period (77 min vs. 62 min, $z=3.32,p<0.001$) and acrophase (68 min vs. 54 min, $z=2.97,p=0.003$) for PWE compared to controls, but not in amplitude (1.98 bpm vs. 2.05 bpm, $z=-0.66,p=0.51$). For PWE, we did not find any correlations between seizure frequency and intra-individual variability in circadian properties, or any difference between weeks with and without seizures.
Interpretation: This finding indicates that the circadian rhythm of heart rate is more variable for people with epilepsy and that this can be detected using a wearable device. However, we were unable to find any associations with seizure frequency or occurrence, suggesting intra-individual variability could be another manifestation of epilepsy aetiology, and not simply reflecting seizure-induced effects.
Putting Research Into Context
Evidence before this study: We searched Scopus, PubMed and Google Scholar up to Nov 20, 2024 using combinations of the following Title/Abstract search terms: circadian AND (rhythm* OR cycle*) AND (disrup* OR variability) AND (epilep* OR seizure*) AND (heart rate OR wearable OR ECG). The majority of the evidence points to seizures occurring at particular phases of the circadian cycle in some patients. However, disruption to the circadian rhythm itself in epilepsy has not been investigated to the same extent, and we found no other study using heart rate measurements to do so in humans.
A broader search for circadian disruptions in general epilepsy using search terms: “circadian AND (rhythm* OR cycle*) AND (disrup* OR day-to-day variability OR intra-individual variability) AND epilep*” revealed much interest in this field, but no description of intra-individual variability of the circadian rhythm.
Added value of this study: We report, for the first time, evidence of long-term circadian disruption in epilepsy captured through increased intra-individual variability of the circadian rhythm of heart rate. Comparing 143 people with epilepsy (PWE) to 31 controls, we found a substantial increase in the intra-individual variability of two out of three circadian rhythm properties. In PWE, we further found no correlation between intra-individual variability and seizure frequency, and no difference in intra-individual variability between weeks with and without seizures.
Implications of all the available evidence: We have established that increased intra-individual variability in the circadian rhythm appears to be a characteristic of epilepsy, independent of seizure occurrence. This implies that the intra-individual variability may be a result of the epilepsy aetiology rather than caused by seizures or their associated consequences. Future work should establish if reducing this variability, e.g. by establishing regular sleep/wake times and scheduled exercise and food intake, could improve patient symptoms and seizures.
@@ 1 Introduction
The circadian rhythm aligns our physiology and behaviour to the 24-hour environmental light-dark cycle. A stable circadian rhythm is thought to be important for overall health, and disruptions to the circadian rhythm have been associated with various conditions, including sleep disorders (1, 2), psychiatric disorders (3, 4, 5, 6), and neurological disorders (7). However, it remains unclear how circadian disruption should be assessed over the long term, with different suggestions across different contexts (8, 9, 10).
One promising and easy-to-interpret approach is to assess the intra-individual variability of some descriptive properties of the circadian rhythm, such as amplitude, measured across multiple days. This approach has also been referred to as ‘day-to-day variability’ (11, 12) or ‘circadian variability’ (4). Intra-individual variability is designed to detect persistent irregularities in circadian rhythms over long-term physiological recordings. Such data are most easily obtained from peripheral measures using wearable devices (such as heart rate, physical activity, and skin temperature), making this approach a cost-effective and scalable methodology to study circadian disruption, complementary with existing methods.
Epilepsy is a neurological disorder characterized by recurrent seizures (13), which are influenced by sleep patterns (14, 15, 16, 17, 18) and the circadian rhythm (19, 20, 21, 22, 23, 24, 25). However, the impact of persistent, long-term circadian disruption on epilepsy remains unstudied. We propose using an intra-individual variability assessment on data from people with epilepsy (PWE), which could provide valuable insights for clinical applications, such as enhancing wearable seizure prediction (26) and guiding chronotherapeutic treatments (20, 25), including ASM alert systems. Additionally, demonstrating the effectiveness of using the intra-individual variability of the circadian rhythm as a measure of disruption in epilepsy could support its broader use as a clinical biomarker in wearable devices.
Here, we measure the intra-individual variability of three descriptive properties of the circadian rhythm of heart rate using long-term wearable-derived heart rate recordings from 143 PWE and 31 controls. We seek to establish whether variability is greater in epilepsy and, if so, whether a relationship exists between variability and seizure frequency and occurrence.
2 Methods
@@ 2.2.3 Computing circadian properties
The extracted CRHR was then split-up into the individual (approximately daily) circadian cycles (Figure 2E), and the period, acrophase and amplitude were computed for each cycle (Figure 2G). To achieve this, the Hilbert transform was applied to the extracted CRHR, producing a complex-valued analytic signal, from which the circadian “phase series” (Supplementary S1.2), a measure of circadian cycle progression, was derived (Figure 2E(ii)). The rhythm was then split into daily cycles at each trough in the phase-series. For each individual cycle, period was computed from the duration between the first and last time point of the cycle, acrophase as the time of day at zero cycle phase, and amplitude as the difference between magnitude at acrophase and trough (Figure 1).
@@ 2.2.4 Calculating the intra-individual variability of circadian properties
Circadian properties for each cycle were grouped into consecutive, non-overlapping segments of seven days. The standard deviation was calculated across all seven days in each segment and property (Figure 2G), and the summary intra-individual variability was calculated from the average of the standard deviations across all segments. The intra-individual average of circadian properties, only used for reference later, was calculated in the same way from the average of the means across segments. The fixed segment size of seven days prevents any bias introduced by varying recording duration and ‘run’ length between participants when calculating the mean and standard deviation (see Supplementary S1.3 for more detail).
@@ 2.3 Statistical Analysis
We calculate p-values for reference only, and all reported values are raw, uncorrected by FDR. A two-sided Wilcoxon rank-sum test was used to compare the intra-individual average and variability of each circadian property between PWE and controls. To account for the imbalance between PWE and control sample sizes, a random sub-sampling test was performed over 10,000 iterations where the whole control sample was compared to a random sub-sample of 28 PWE using the same statistical method as above. Sampling was without replacement within an iteration, but with replacement across iterations.
@@
To test whether having one or more seizures in a 7-day period was associated with greater intra-individual variability, each 7-day segment occurring within a participant’s seizure diary was classified as either seizure-containing or seizure-free according to the seizure diary. For PWE with at least 5 seizure-free segments and at least 5 seizure-containing segments, the intra-individual circadian average and variability was calculated separately over seizure-free and seizure-containing 7-day segments. These values were then compared using a two-sided Wilcoxon signed-rank paired test.
2.4 Role of the funding source
Funders played no role in study design; nor in the collection, analysis, and interpretation of data; nor in the writing of the report; nor in the decision to submit the paper for publication.
@@ 3 Results
143 people with epilepsy (PWE) and 31 controls remained after processing. Supplementary table 1 lists the demographics and other metadata of this cohort. An average of 163 days (sd = 183) were discarded for each subject, resulting in an average recording duration of 443.9 days (sd = 530.9) for PWE and 221.7 days (sd = 257.0) for controls. Although recording duration was shorter in controls compared to PWE, we found no relationship between (post-segmenting) recording duration and intra-individual variability in any circadian property (Supplementary S3). There were no age ($t=-0.215,p=0.84$) or sex ($\chi^{2}=2.777,p=0.427$) differences between PWE and controls, and there was no association between intra-individual variability and age, sex or epilepsy sub-type (see Supplementary S5).
@@ 3.1 The circadian rhythm of heart rate is more variable for people with epilepsy
The intra-individual variability in period (77min vs. 62 min, $z=3.32,p<0.001$, Figure 3A) and acrophase (68 min vs. 54 min, $z=2.97,p<0.001$, Figure 3B) was increased for PWE compared to controls. However, there was no difference in intra-individual variability of amplitude between PWE and controls ($\sim$2bpm, $z=-0.66,p=0.51$, Figure 3C).
For reference, we also present the results for intra-individual average in Figure 3, but found no substantial difference between PWE and controls (period: $z=0.36,p=0.72$; acrophase: $z=0.95,p=0.34$; amplitude: $z=-1.52,p=0.13$).
All results held after applying random sub-sampling to equal numbers of PWE and controls (see Supplementary S4).
Figure 3: Comparison of the distribution of the intra-individual average and variability of circadian properties between PWE and controls. Period, acrophase and amplitude are defined in Figure 1. The two-sided Wilcoxon rank-sum test was used for comparison between PWE and Controls. Each subpanel has been zoomed-in to omit outliers; outliers were not excluded prior to statistical testing.
3.2 Intra-individual variability of circadian properties does not correlate with increased seizure frequency
Next, the correlation between seizure frequency and intra-individual variability was investigated (Figure 4). PWE without any seizures recorded or with outlying intra-individual variability (absolute z-score $>$ 3) in any property were removed, leaving 117 PWE. We found that intra-individual variability of any property was not correlated with seizure frequency ($R^{2}<<0.1$).
Figure 4: Scatter plots between intra-individual variability of circadian properties and seizure frequency. Seizure frequency is in units of seizures per week on log10 scale. n=117 for all panels.
3.3 The intra-individual variability of circadian properties does not differ between weeks with and without seizures
Finally, we investigated whether intra-individual variability differed between seizure-containing or seizure-free 7-day segments (Figure 5). Individuals without at least 5 seizure-free segments and at least 5 seizure-containing segments were excluded, leaving 57 PWE. There was no paired difference in intra-individual variability of any circadian property between seizure-free vs. seizure-containing segments. However, there was a decrease in the intra-individual average of amplitude in weeks with seizures ($z=-2.15,p=0.03$, Figure 3C), possibly indicating a change in behaviour (being less active during periods with seizures). However, there was no difference for average period or acrophase.
@@ 4 Discussion
This study analysed the intra-individual variability in three properties of the circadian rhythm using long-term wearable heart rate data from 143 PWE and 31 controls. We found an increased variability in circadian period and acrophase for PWE, providing initial evidence of circadian disruption in epilepsy. However, we found no evidence of an association between intra-individual variability and seizure frequency, or occurrence at the population level.
@@
Circadian disruption has a bi-directional relationship with health (27) and is associated with negative health outcomes (10) and mortality over the long term (28). It is therefore likely to have compounding health implications for PWE, so the results presented here should motivate the development of interventions to reduce the co-morbidities associated with circadian disruption in PWE (29).
@@
We found no evidence of any link between disruption and seizures but consider a link with epilepsy aetiology possible. Evidence from ex vivo tissue (30), and intracranial EEG (22) indicates that tissue associated with epilepsy pathology may have an impaired circadian ‘clock’, which could also have downstream effects on the expression of the circadian rhythm in heart rate. Co-morbidities of epilepsy such as anxiety and depression (31, 32) are also themselves associated with circadian disruption (33), so may also contribute to the observed difference. Within the epilepsy cohort we found that intra-individual variability has a wide range and considerable overlap with the control distribution: future work should investigate the extent to which the co-morbidities mentioned above can explain these findings.
Misalignment of the circadian rhythm with the environmental light/dark cycle, particularly towards the evening, is one form of circadian disruption that has been associated with psychiatric disorders (6, 10). Chronotype studies in epilepsy have proven less conclusive (25). Precisely reporting on chronotype is outside the scope of our study - average acrophase most closely resembles chronotype, though no differences were observed between PWE and controls, or between focal and generalised sub-types. However, increased acrophase variability for PWE may be associated with circadian misalignment.
@@
Our approach has some key limitations. Firstly, while heart rate follows a strong circadian cycle (34), bouts of activity such as exercise or stress can cause ‘behavioural masking’, interfering with measurement of the circadian state (35), particularly when derived from wearable PPG. Despite the issues of noise, resolution and accuracy associated with wearable devices, they provide key advantages in affordability, convenience and scalability over traditional methods for measuring the circadian rhythm such as melatonin sampling, enabling their wider use in clinical contexts and normative benchmarking.Secondly, our analysis does not include some key variables that may play a role in the intra-individual variability we observe. Anti-seizure medications (ASMs) have been reported to both stabilise and disrupt sleep (15, 16), and so could explain the increased intra-individual variability we see in PWE as well as the overlap with controls. Demographics such as age and sex are also known to associate with circadian parameters (7, 36). We did observe a correlation between age and intra-individual averages of acrophase and amplitude as expected based on this previous literature, but not with variability in any property. We did not find any sex differences with respect to circadian average or variability. Future work should take advantage of multi-modal wearables, integrating heart rate with activity and temperature for circadian state estimation, for example.
@@
y, the role of sleep disruption in epilepsy is well noted (37, 21, 38, 16, 39, 17, 20), and while a distinct process physiologically (21, 40, 8), it is intrinsically tied to the circadian rhythm. As the circadian behaviour of heart rate is tied to sleep, its variability may primarily driven by sleep disruptions. Future studies should incorporate sleep quality questionnaires and wearable sleep tracking to distinguish between sleep-related variability vs. more broad intra-individual variability of the circadian rhythm.

### v4

# More variable circadian rhythms in epilepsy captured by long-term heart rate recordings from wearable sensors

@@ Abstract
Objective: The circadian rhythm synchronizes physiological and behavioural patterns with the 24-hour light-dark cycle. Disruption to the circadian rhythm is linked to various health conditions, though optimal methods to describe these disruptions remain unclear. An emerging approach is to examine the intra-individual variability in measurable properties of the circadian rhythm over extended periods. Epileptic seizures are modulated by circadian rhythms, but the relevance of circadian rhythm disruption in epilepsy remains unexplored. Our study investigates intra-individual circadian variability in epilepsy and its relationship with seizures.
Methods: We retrospectively analyzed over 70,000 hours of wearable smartwatch data (Fitbit) from 143 people with epilepsy (PWE) and 31 healthy controls. Circadian oscillations in heart rate time series were extracted, daily estimates of circadian period, acrophase, and amplitude properties were produced, and estimates of the intra-individual variability of these properties over an entire recording were calculated.
Results: PWE exhibited greater intra-individual variability in period (76 min vs. 57 min, $d=0.66,p<0.001$) and acrophase (64 min vs. 48 min, $d=0.49,p=0.004$) compared to controls, but not in amplitude (2 bpm, $d=-0.15,p=0.49$). Variability in circadian properties showed no correlation with seizure frequency, nor any differences between weeks with and without seizures.
Significance: For the first time, we show that heart rate circadian rhythms are more variable in PWE, detectable via consumer wearable devices. However, no association with seizure frequency or occurrence was found, suggesting that this variability might be underpinned by the epilepsy aetiology rather than being a seizure-driven effect.
Key points
  • Relative to controls, the circadian rhythm of heart rate was more variable for PWE, measured using commercial wearable devices.
  • This provides evidence for long-term and persistent circadian disruption in epilepsy.
  • However, increased intra-individual circadian variability was not related to seizure occurrence or frequency.
  • The underlying aetiology may instead be associated, or factors such as ASMs or comorbidities may play an important role.
@@ 1 Introduction
The circadian rhythm aligns our physiology and behaviour to the 24-hour environmental light-dark cycle. A stable circadian rhythm is thought to be important for overall health, and disruptions to the circadian rhythm have been associated with various conditions, including sleep disorders (1, 2), psychiatric disorders (3, 4, 5, 6), and neurological disorders (7). However, it remains unclear how circadian disruption should be assessed over the long term - many methods are available depending upon context (8, 9, 10).
One promising and easy-to-interpret approach is to assess the intra-individual variability of a set of fundamental descriptive properties of the circadian rhythm, for example amplitude, measured across multiple consecutive days. This approach has also been referred to as ‘day-to-day variability’ (11, 12) or ‘circadian variability’ (4). Intra-individual variability is designed to detect persistent irregularities in circadian rhythms over long-term physiological recordings. Such data are most easily obtained from peripheral measures that exhibit circadian oscillations using wearable devices (such as heart rate, physical activity, and skin temperature), making this approach a cost-effective and scalable methodology to study circadian disruption, complementary with existing methods.
Epilepsy is a neurological disorder characterized by recurrent seizures (13), which are closely associated with the circadian rhythm (14, 15, 16, 17, 18, 19, 20). Sleep patterns are inherently tied to the circadian rhythm, and the relationship between sleep and seizures has been reported upon widely (21, 22, 23, 24, 25). However, the role of persistent, long-term circadian disruption in epilepsy remains unstudied. Consequently, we propose the assessment of intra-individual circadian variability in wearable heart rate recordings sourced from people with epilepsy (PWE), which could provide valuable insights for clinical applications, such as enhancing wearable seizure prediction (26) and guiding chronotherapeutic treatments (15, 20), including ASM alert systems. Additionally, this novel application of intra-individual variability as a measure of circadian disruption in epilepsy could support its broader use as a wearable clinical biomarker.
Here, we measure the intra-individual variability of three descriptive properties of the circadian rhythm of heart rate using long-term wearable-derived heart rate recordings from 143 PWE and 31 controls. We seek to establish whether circadian variability is greater in epilepsy and, if so, whether a relationship exists between variability and seizure frequency and occurrence.
2 Materials & methods
@@ 2.2.3 Computing circadian properties
The extracted CRHR was then split-up into the individual (approximately daily) circadian cycles (Figure 2E), and the period, acrophase and amplitude were computed for each cycle (Figure 2G). To achieve this, the Hilbert transform was applied to the extracted CRHR, producing a complex-valued analytic signal, from which the circadian “phase series” (Supplementary S1.2), a measure of circadian cycle progression, was derived (Figure 2E(ii)). The rhythm was then split into daily cycles at each trough in the phase-series. For each individual cycle, period was computed as the duration between the first and last time point of the cycle, acrophase as the time of day at zero cycle phase, and amplitude as the difference in magnitude between acrophase and trough (Figure 1).
@@ 2.2.4 Calculating the intra-individual variability of circadian properties
Circadian properties for each cycle were grouped into consecutive, non-overlapping segments of seven days. The standard deviation was calculated using all seven days in each segment and property (Figure 2G), and the summary intra-individual variability was calculated from the average of the standard deviations across all segments. The intra-individual average of circadian properties was calculated in the same way from the average of the means across segments. The fixed segment size of seven days prevents any bias introduced by varying recording duration and ‘run’ length between participants when calculating the mean and standard deviation (see Supplementary S1.3 for more detail).
@@ 2.3 Statistical analysis
We calculate p-values for reference only, and all reported values are raw, uncorrected by FDR. A non-parametric two-sided Wilcoxon rank-sum test was used to compare the intra-individual average and variability of each circadian property between PWE and controls. A non-parametric test was chosen to account for the imbalance between PWE and control sample sizes, but to further control for this a random sub-sampling test was performed over 10,000 iterations, where the whole control sample was compared to a random sub-sample of 28 PWE using the same statistical method as above. Sampling was without replacement within an iteration, but with replacement across iterations.
@@
To test whether having one or more seizures in a 7-day period was associated with greater intra-individual variability, each 7-day segment occurring within a participant’s seizure diary was classified as either seizure-containing or seizure-free according to the seizure diary. For PWE with at least 5 seizure-free segments and at least 5 seizure-containing segments, intra-individual circadian average and variability values were calculated over seizure-free and seizure-containing segments separately. These values were then compared using a two-sided Wilcoxon signed-rank paired test.
Effect sizes were reported using Cohen’s d throughout.
@@ 3 Results
Before processing, data were available for n=164 people with epilepsy (PWE) and n=36 controls. After processing, 143 PWE and 31 controls remained for analysis. We analysed and compared intra-individual variability of circadian rhythm properties (see Methods) calculated using wearable heart rate data (FitBit), between PWE and controls.
Briefly, we isolated the circadian rhythm component from the raw heart rate recordings, and measured a range of properties (period, acrophase, amplitude) for each cycle. Intra-individual variability in circadian properties were then obtained across segments of successive cycles. An increased intra-individual variability in any property can be interpreted as a less “stable” circadian rhythm for the individual.
Supplementary table 1 lists the demographics and other metadata of this cohort. An average of 163 days (sd = 183) were discarded for each subject, resulting in a median recording duration of 215.8 days (IQR = 580.9) for PWE and 125.1 days (IQR = 260.5) for controls. Although recording duration was shorter in controls compared to PWE, we found no relationship between (post-segmenting) recording duration and intra-individual variability in any circadian property (Supplementary S3). There were no age ($t=-0.215,p=0.84$) or sex ($\chi^{2}=2.777,p=0.427$) differences between PWE and controls, and there was no association between intra-individual variability and age, sex or epilepsy sub-type (see Supplementary S5).
@@ 3.1 The circadian rhythm of heart rate is more variable for people with epilepsy
The intra-individual variability in period (76 min vs. 57 min, $d=0.66,p<0.001$, Figure 3A) and acrophase (64 min vs. 48 min, $d=0.49,p=0.004$, Figure 3B) was increased for PWE compared to controls. However, there was no difference in intra-individual variability of amplitude between PWE and controls ($\sim$2 bpm, $d=-0.15,p=0.49$, Figure 3C).
For reference, we also present the results for intra-individual average in Figure 3, but found no substantial difference between PWE and controls (period: $d=-0.17,p=0.66$; acrophase: $d=0.04,p=0.35$; amplitude: $d=-0.22,p=0.13$).
All results held after applying a repeated sub-sampling test with randomly selected samples of PWE equal in size to controls (see Supplementary S4).
Figure 3: Comparison of the distribution of the intra-individual average and variability of circadian properties between PWE and controls. Period, acrophase and amplitude are defined in Figure 1. Effect sizes are measured in Cohen’s d and p-values are based on the two-sided Wilcoxon rank-sum test for comparison between PWE and Controls. Each subpanel has been zoomed-in to omit some outliers; outliers were not excluded for statistical testing.
3.2 Intra-individual variability does not correlate with increased seizure frequency
Next, the correlation between seizure frequency and intra-individual variability was investigated (Figure 4). PWE without any seizures recorded or with outlying intra-individual variability (absolute z-score $>$ 3) in any property were removed, leaving 119 PWE. We found that intra-individual variability of any property was not correlated with seizure frequency ($|r|\leq 0.1$).
Figure 4: Scatter plots between intra-individual variability of circadian properties and seizure frequency. Seizure frequency is in units of seizures per week on log10 scale.
3.3 Intra-individual variability does not differ between weeks with and without seizures
We investigated whether intra-individual variability differed between seizure-containing or seizure-free 7-day segments (Figure 5). Individuals without at least 5 seizure-free segments and at least 5 seizure-containing segments were excluded, leaving 56 PWE. There was no paired difference in intra-individual variability or average of any circadian property between seizure-free vs. seizure-containing segments.
@@ 3.3 Intra-individual variability does not differ between weeks with and without seizures
3.4 Weekend and time of year effects
To supplement our findings, we investigated (Supplementary S6) whether weekdays and weekends, as well as the time of year were associated with the observed increased intra-individual variability in PWE.
We observed a difference between weekends and weekdays in almost all measures (intra-individual averages and variabilities) for both controls and PWE (Fig S7). However, the previously observed increases in intra-individual variability in period and acrophase for PWE remained regardless of weekday or weekend (Fig. S8). This observation was additionally confirmed by a mixed effects model, where we tested for the interaction of weekday/weekend with control/PWE. More details and full statistical models can be found in Suppl. S6.1.
We also tested for a seasonal effect in a similar manner, where we found a less pronounced effect overall. We again confirmed in that the observed increase in intra-individual variability on period and acrophase were not driven by seasonal effects or interactions. More details and full statistical models can be found in Suppl. S6.2.
In summary, whilst weekly and seasonal effects are present in intra-individual averages and variability, these do not drive our main results of increased intra-individual variability in PWE.
@@ 4 Discussion
This study analysed the intra-individual variability of three properties of the circadian rhythm using long-term wearable heart rate data from 143 PWE and 31 controls. We found increased variability in circadian period and acrophase for PWE, providing crucial evidence of circadian disruption in epilepsy. However, we found no evidence of an association between intra-individual variability and seizure frequency, or occurrence, at the population level.
@@
Circadian disruption has a bi-directional relationship with health (27) and is associated with negative health outcomes (10) and mortality over the long term (28). It is therefore likely to have compounding health implications for PWE, so the results presented here should motivate the development of interventions to reduce the co-morbidities associated with circadian disruption for PWE (29).
@@
We found no evidence of any link between circadian disruption and seizures but consider a link with epilepsy aetiology possible. Epilepsy often has a genetic cause (13) and many related genes have been identified. Associations between epilepsy and abnormal expression of the genes underlying the circadian rhythm have also been reported (30, 31). For example, evidence from human ex vivo pathological tissue suggests a link between decreased expression of BMAL1 and temporal lobe epilepsy (32) - similar findings have been reported for the CLOCK gene (33). Functionally, the circadian amplitude of neural activity has been found to be reduced in pathological areas using intracranial EEG (17). It is possible that any circadian rhythm disruption with a genetic basis at the brain level could be reciprocated at the heart, either genetically or as a downstream effect of disruption to the central circadian rhythm - such effects may explain the disruption to the circadian rhythm of heart rate for people with epilepsy that we observe here. Moreover, co-morbidities of epilepsy such as anxiety and depression (34, 35) are also associated with circadian disruption (36). Within the epilepsy cohort we found that intra-individual variability has a wide range and considerable overlap with the control distribution: future work should investigate the extent to which these co-morbidities can explain our findings.
Misalignment of the circadian rhythm with the environmental light/dark cycle, particularly towards the evening, is one form of circadian disruption that has been associated with psychiatric disorders (6, 10). ‘Chronotype’ - an individual’s innate preference as a morning or evening type - relates to such misalignment, and has been investigated in epilepsy (37, 30, 20). Precisely reporting on chronotype requires specific methodology so is outside the scope of our study, though average acrophase most closely resembles chronotype. No differences in acrophase were observed between PWE and controls, or between focal and generalised sub-types in our data (Supplementary S6). Incidentally, we did observe some differences in circadian rhythms between weekends and weekdays (Figure S7) for both PWE and controls. In particular the differences in intra-individual average period and acrophase may be related to the ‘social jetlag’ phenomenon, which refers to differences in sleep and behavioural patterns on weekdays compared to weekends due to the constraints of work, education or other commitments (8). These changes in intra-individual averages were different for PWE compared to controls, which may support previous findings that social jetlag is exacerbated for PWE (37). However, most weekday vs. weekend effects in intra-individual variability were independent of the control vs. PWE effects (Figure S8, Table 3). Specifically, the highlighted increased intra-individual variability in PWE was not driven by the weekday/weekend effect in our data. It should be noted that insights into social jetlag are limited without inclusion of sleep measures in this study. Future work could test, in a more formal way, the mechanistic relationship between social jetlag, chronotypes, sleep, and our proposed measures of circadian variability.
Interestingly, in this study we observed that the median circadian period was marginally longer than 24 hours (24.02h) for both controls and PWE. In constant laboratory conditions without time cues such as light/dark cycles, the circadian rhythm persists in humans, but with a period of up to 25 hours (38). In a laboratory setting where light/dark cycles are fixed, (39) report a variable circadian period of temperature and melatonin rhythms between participants, above and below 24 hours, that was shorter (24.09h) on average for females than males (24.19h). Our study was undertaken under no such controlled conditions: circadian period would certainly be entrained by light and other environmental cues, which may explain why our reported average is closer to 24 hours. Additionally, we did not observe any difference in average circadian period between males and females (Supplementary S5). Shifts in average circadian period may be associated with modulatory effects of longer, multi-day rhythms of heart rate (14), or associated with the annual variations of circadian properties that we also observed (Section S6.2). Future work could explore the dynamics of circadian period over-time in detail and whether re-entrainment differs between PWE and controls.
@@
Our approach has some key limitations. Firstly, while heart rate follows a strong circadian cycle (40), bouts of activity such as exercise or stress can cause ‘behavioural masking’, interfering with measurement of the circadian state (41), particularly when derived from wearable PPG. Despite the issues of noise, resolution and accuracy associated with wearable devices, they provide key advantages in affordability, convenience and scalability over traditional methods for measuring the circadian rhythm such as melatonin sampling, enabling their wider use in clinical contexts and normative benchmarking. Future work should take advantage of multi-modal wearables, integrating heart rate with activity and temperature for circadian state estimation, for example. Secondly, our analysis does not include some key variables that may play a role in the intra-individual variability we observe. Anti-seizure medications (ASMs) have been reported to both stabilise and disrupt sleep (22, 23), and so could explain the increased intra-individual variability we see in PWE as well as the overlap with controls if this effect translates to the circadian rhythm more generally. Demographics such as age and sex are also known to associate with measures of the circadian rhythm (7, 42). In a subset of the cohort (Supplementary S5), we did observe a correlation between age and intra-individual averages of acrophase and amplitude as expected based on this previous literature, but not with variability in any property. We did not find any sex differences with respect to circadian average or variability. Future work should further clarify whether associations with these potential co-variates can be ruled out.
@@
Finally, the role of sleep disruption in epilepsy is well noted (43, 16, 44, 23, 45, 24, 15), and while a distinct process physiologically (16, 46, 8), it is intrinsically tied to the circadian rhythm. As the circadian behaviour of heart rate is tied to sleep, its variability may primarily be driven by sleep disruptions. Future studies should incorporate sleep quality questionnaires and wearable sleep tracking to distinguish between sleep-related variability vs. more broad intra-individual variability of the circadian rhythm.
5 Conclusion